shift summary Evening 5/26/99

Herman B. White (hwhite@fnal.gov)
Thu, 27 May 1999 00:13:36 -0500 (CDT)

Shift Summary for Evening shift, 05/26/99

On Shift: Jason Hamm
Herman White
A variety of other committed collaborators were working throughout this
shift.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 8.5 hrs

Accounting of lost time:

No beam this shift due in part to preparations for operations and
instrumentation installation. Electronics work is also in progress on the
detector.

HV mainframe #14 is turned off, with possible faulty cards. Rick Tesarek
expects to investigate tomorrow, and evaluate the condition of the HA HV
and bases.

Sydney T. made modifications to the DIS data base, which will take effect
after run 13144. These modifications are necessary to accommodate the new
15Gb DLT III XT tapes. see details Log XVIII p.300

Crate 57 was replaced again. The cause of this crate failure was not
noted.

D. Jensen investigated problems with the z-calibrator. There is some "air
puffing" detected from a valve when the z-calibrator is in the down
position. Investigation continues. Vacuum measurements indicate a level at
~1-1.8 E-06 (torr).

Doug also initiated measurements of the z-positions of the Regen, DC1, DC2
with the laser-ranger equipment.

(JAWS & EdB) Work continues on the calorimeter. HV & DPMT channels
(1817/1818) were swapped, to address wide RMS in laser stability. No definitive
conclusions that this resolves the problem. Experts suggests that channel
1817 be watched for stability problems.
The HV divider for CsI 440 was swapped, to illiminate missing signal in
the laser illumination plot.

(S. Glazov and Miguel B.) fixed crate #24 readout for drift chamber x-view
readout problems. A problem still persist running this crate at high
rate.

Arun Tripathi installed a new version of KT-Status on the online machine.

========================================================================
Runs ended this shift:
Run 13146 - completed
Run 13147 - cosmic run (still running at the end of the shift)

Tapes brought to FCC this shift: 0